a coin is weighted so that there is a 60.7% chance of it landing on heads when flipped. the coin is flipped 15 times
find the expected number of flips that will result in heads
round answer 2 decimal places
Answer #1

Solution :

Given that,

p = 0.607

q = 1 - p = 1 - 0.607 = 0.393

n = 15

Using binomial distribution,

expected number = = n * p = 15 * 0.607 = 9.11